THE SUBMILLIMETER-WAVE SPECTRUM OF THE CrH AND CrD (^{6}Σ^{+}) RADICALS

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The N = 0 → 1 transition of CrH and the N = 1 → 2 line of CrD in their ^{6}Σ^{+} ground states have been recorded using sub-millimeter direct absorption techniques. This study was the first direct measurement of these transitions. These molecules were produced in a discharge plasma of chromium vapor and H_{2} or D_{2} gas. Three fine structure components were measured for each radical, and the proton and deuteron hyperfine splittings were also resolved. Spectroscopic constants were obtained by fitting the data with a Hund’s case (b) Hamiltonian including rotation, spin-rotation, spin-spin, magnetic hyperfine, and electric quadrupole interactions. These measurements have resulted in improved spectroscopic constants for both isotopomers. The hyperfine parameters suggest that the bonding in CrH has a nonnegligible covalent character. This molecule may be present in the outer envelope of late-type stars and formed by ion-molecule reactions from neutral or ionized Cr and hydrogen gas. Chromium has a cosmic abundance of Cr/H = 5 x 10^{−7}.